OpenStudy (anonymous):

what are the intervals of y=-2tan2/3x?

OpenStudy (anonymous):

I assume this is the function: \[y=-2\tan \left( \frac{ 2 }{ 3 } x \right) \]

OpenStudy (anonymous):

For the tangent function you have to make restrictions. Whenever I work with tangent I just recall that: \[\tan =\frac{ \sin }{ \cos } \] This straight away tells me the restriction. As you CANNOT divide by 0. Cos function can take 0s, we do not want that to happen.

OpenStudy (anonymous):

Let us just solve this first: cos x= 0 That is at \[x=\frac{ \pi }{ 2 }\] However that is not the full solution as all the trigonometric functions are PERIODIC! Thus the correct solution is \[x=\frac{ \pi }{ 2 } +2k \pi \] where k is any integer, \[k \in \mathbb{Z} \]

OpenStudy (anonymous):

But I solved cosx=0 and for this question we need: cos(2/3 x)=0 The above helps us greatly: \[\frac{ 2 }{ 3 }x =\frac{ \pi }{ 2 } +2k \pi \] Now multiply through by 3/2 to get only x on the left hand side \[x=\frac{ 3\pi }{ 4 }+3k \pi \] where \[k \in \mathbb{Z}\] This is where your function does not make sense, at any other points it is perfectly fine. The range of tangent is always every number \[R \in \mathbb{R} \]